OpenStudy (anonymous):

Please help with this algebra question! Medals will be given. Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (anonymous):

that means start with 1, 2 then add the last two terms to get the next one \[1,2,2+1=3,3+2=5,5+3=8,...\]

OpenStudy (anonymous):

i.e \[1,2,3,5,8\]

OpenStudy (ellecullen):

f(3) = f(3-1) + f(3-2) 3>2 f(3) = f(2)+f(1) like 2+1 = 3 try this tip an let me know if it works elle